OpenStudy (anonymous):

a sheet of paper is 8 1/2 in wide and 11 in long. the sheet is printed with a margin of 1 1/4 in wide on all four sides. Find the perimeter of the printed side of the page.

OpenStudy (anonymous):

nvm answer 29. 8.5 - (1.25x2) = 6 x2 = 12 11 - (1.25x2) = 8.5 x2 = 17 12+ 17 =29 inches